# Natural polynomials for Kerr quasi-normal modes

### v2

@@ VII Select numerical results
In Eq. (100), $u_{k}^{(a)}$ are canonical confluent Heun polynomials derived from a BH spin of $a$, where $(s,l,m,n)$ are implicitly fixed. Vectors $\vec{f}_{p}^{(0)}$ and $\vec{f}_{q}^{(a)}$ will be considered normalized, making e.g. the set of all $\vec{f}_{p}^{(0)}$ orthonormal. Therefore, the decomposition of a Kerr radial function onto the basis of Schwarzschild radial functions is
$\vec{f}_{q}^{(a)}\;=\;\sum_{q=0}\;\vec{f}_{p}^{(0)}\;(\vec{f}_{p}^{(0)}\cdot\vec{f}_{q}^{(a)})\;.$ (101)
In Eq. (101), $(\vec{f}_{p}^{(0)}\cdot\vec{f}_{q}^{(a)})$ are elements of the mixing matrix between the Schwarzschild and Kerr radial functions. In the case of $a=0$, Eq. (101) simply reduces to $\vec{f}_{q}^{(a)}=\sum_{q=0}\vec{f}_{p}^{(0)}\delta_{pq}$, signaling a diagonal mixing matrix.
In Fig. 7, the $(s,\ell,m,n)=(-2,2,2,0)$ QNM frequencies have been used to compute Schwarzschild-Kerr mixing matrices for four values of BH spin666 The successful reproduction of Fig. 7 requires that one keep in mind the overall sign ambiguity discussed in Sec. V: Given canonical polynomials, $u_{p}^{(a)}$ and $u_{p}^{(a^{\prime})}$, one can avoid the presence of an incongruous minus sign by using, e.g., the scalar product space of $u_{p}^{(a)}$ to compute the matrix whose diagonal elements are ${\langle{u_{p}^{(a)}}\,|\,{u_{p}^{(a^{\prime})}}\rangle}$. Diagonal elements with a negative real part signal that $u_{p}^{(a^{\prime})}$ should be scaled by $-1$. . As seen from left to right in Fig. 7, as BH spin increases, so too does the magnitude of off-diagonal matrix elements. Nevertheless, the mixing matrix appears to be approximately diagonal throughout the range of spins shown.
@@ VIII Discussion and conclusions
The present work is based upon a simple premise: to solve Teukolsky’s radial equation, rather than looking to special functions from other problems, we need only look to the equation itself London (2023a). The result is a method for computing global solutions to the confluent Heun equation via a single expansion in polynomials that are natural to the problem. We have introduced such polynomials in Sec. V. In Sec. VI, we have shown that these polynomials enable the radial equation to be represented as a matrix that is both tridiagonal and symmetric. Lastly, in Sec. VII, we have demonstrated the use of these polynomials in various numerical examples, the last of which may be of particular relevance to the time-dependent perturbations of BHs Green et al. (2023); Cannizzaro et al. (2023); Redondo-Yuste et al. (2023); Sberna et al. (2022); Lagos and Hui (2022). We now briefly discuss connections and implications that have not been fully articulated thus far.
@@
Of the techniques presented here, a number are open to refinement, further exploration, and connection with other recent results. For example, we have shown that, in some cases, the QNM’s separation constant is well approximated by the eigenvalue of a confluent Heun polynomial. We expect that more analytic understanding is tractable when either representing the problem in confluent Heun polynomials, or their canonical counterparts. Further, there is good reason to think that more may be learned by a thorough comparison between our results, and those known for Pollaczek-Jacobi polynomials Chen and Dai (2010); Chen et al. (2019); Min and Chen (2023); Yu et al. (2022). This effort is ongoing and near completion Gurevich and London (2024). There is also the possibility of combining the results presented here with recent iterative generalized spectral methods for computing QNM frequencies Chung et al. (2023a); Blázquez-Salcedo et al. (2023); Chung et al. (2023b); Ghojogh et al. (2023); Zhu et al. (2023).

### v3

@@ VII Select numerical results
In Eq. (100), $u_{k}^{(a)}$ are canonical confluent Heun polynomials derived from a BH spin of $a$, where $(s,l,m,n)$ are implicitly fixed. Further, bra-ket operations use the weight defined by the labeled spin value (e.g. $\langle{f_{p}^{(0)}}\,|\,{u_{k}^{(0)}}\rangle$} uses Eq. (14) defined with $a=0$). Vectors $\vec{f}_{p}^{(0)}$ and $\vec{f}_{q}^{(a)}$ will be considered normalized, making e.g. the set of all $\vec{f}_{p}^{(0)}$ orthonormal. Therefore, the truncated set of Schwarzschild radial functions may constitute an orthonormal basis with a well defined projection operator. The application of this projection operator to $\vec{f}_{q}^{(a)}$ is equivalent to the decomposition of a Kerr radial function onto the basis of Schwarzschild radial functions. This is
$\vec{f}_{q}^{(a)}\;=\;\sum_{p=0}\;\vec{f}_{p}^{(0)}\;(\vec{f}_{p}^{(0)}\cdot\vec{f}_{q}^{(a)})\;.$ (101)
In Eq. (101), $\vec{f}_{p}^{(0)}\cdot\vec{f}_{q}^{(a)}$ are elements of the mixing matrix between the Schwarzschild and Kerr radial functions. In the case of $a=0$, Eq. (101) simply reduces to $\vec{f}_{q}^{(a)}=\sum_{q=0}\vec{f}_{p}^{(0)}\delta_{pq}$, signaling a diagonal mixing matrix.
In Fig. 7, the $(s,\ell,m,n)=(-2,2,2,0)$ QNM frequencies have been used to compute Schwarzschild-Kerr mixing matrices for four values of BH spin [^13] The successful reproduction of Fig. 7 requires that one keep in mind the overall sign ambiguity discussed in Sec. V: Given canonical polynomials, $u_{p}^{(a)}$ and $u_{p}^{(a^{\prime})}$, one can avoid the presence of an incongruous minus sign by using, e.g., the scalar product space of $u_{p}^{(a)}$ to compute the matrix whose diagonal elements are $\langle{u_{p}^{(a)}}\,|\,{u_{p}^{(a^{\prime})}}\rangle$}. Diagonal elements with a negative real part signal that $u_{p}^{(a^{\prime})}$ should be scaled by 1-1. . As seen from left to right in Fig. 7, as BH spin increases, so too does the magnitude of off-diagonal matrix elements. Nevertheless, the mixing matrix appears to be approximately diagonal throughout the range of spins shown.
@@ VIII Discussion and conclusions
The present work is based upon a simple premise: to solve Teukolsky’s radial equation, rather than looking to special functions from other problems, we may need only look to the equation itself London (2023a); Fiziev (2009b). The result is a method for computing global solutions to the confluent Heun equation via a single expansion in polynomials that are natural to the problem. We have introduced such polynomials in Sec. V. In Sec. VI, we have shown that these polynomials enable the radial equation to be represented as a matrix that is both tridiagonal and symmetric. Lastly, in Sec. VII, we have demonstrated the use of these polynomials in various numerical examples, the last of which may be of particular relevance to the time-dependent perturbations of BHs Green et al. (2023); Cannizzaro et al. (2023); Redondo-Yuste et al. (2023); Sberna et al. (2022); Lagos and Hui (2022). We now briefly discuss connections and implications that have not been fully articulated thus far.
@@
Of the techniques presented here, a number are open to refinement, further exploration, and connection with other recent results. For example, we have shown that, in some cases (particularly for large $n$), the QNM’s separation constant is well approximated by the eigenvalue of a confluent Heun polynomial. We expect that more analytic understanding is tractable when either representing the problem in confluent Heun polynomials, or their canonical counterparts. Further, there is good reason to think that more may be learned by a thorough comparison between our results, and those known for Pollaczek-Jacobi polynomials Chen and Dai (2010); Chen et al. (2019); Min and Chen (2023); Yu et al. (2022). There is also the possibility of combining the results presented here with recent iterative generalized spectral methods for computing QNM frequencies Chung et al. (2023a); Blázquez-Salcedo et al. (2023); Chung et al. (2023b); Ghojogh et al. (2023); Zhu et al. (2023). Regading spectral pollution, it may also be possible to adapt techinques deveoped in quatum meechanics to mitigate the effect without incuring the added computational cost of increased basis size Lewin and Séré (2008); Boulton (2016); Davies and Plum (2003); Martínez Adame (2007).